# On the Oscillations in Cournot Games with Best Response Strategies

### v2

@@ 1 Introduction
  • We show that the oscillation period cannot exceed two periods. The main challenge to analyze the best response dynamics is due to the non-linearity of the best response strategy. By choosing the greater value between a linear term and zero as the best response strategy, it becomes difficult to derive a straightforward formula for the product quantity in future rounds.
  • We classify all two-period oscillations into three types. The first type occurs when each firm has zero production in one of the two periods. In the second type, fewer than three firms have positive production in the first period, and some additional firms have positive production in the second period. The third type occurs in the case of exact three firms.
  • We propose linear-time algorithms to identify all three types of two-period oscillations.

### v3

@@ 1 Introduction
  • Under linear inverse demand with non-negative quantity constraints, we prove that simultaneous best-response dynamics in an $n$-firm Cournot game must converge exclusively to either a Nash equilibrium or a two-period cycle.
  • We further establish that any two-period cycle must conform to one of three specific structural patterns. The first type arises when each firm produces zero output in one of the two periods. The second type occurs when fewer than three firms produce a positive quantity in the first period, with additional firms producing positive quantities in the second period. The third type arises in the case of exactly three firms.
  • We propose linear-time $O(n)$ algorithms that simultaneously detect oscillation patterns and classify their structural type. This computational contribution equips practitioners with a diagnostic toolkit for empirically analyzing strategic cycling behavior in real-world oligopolistic markets.